Question: The manager of the Inappropriate Compensation Merchandising Co. receives as his or her compensation for the year $700,000 multiplied by the firm's return on assets. Return on assets for any year is defined as the earnings for the year divided by the total assets at the end of the year. If the manager takes no other actions this year, earnings will be $150,000 and total assets at the end of the year will be $1,500,000. This means that, with no further actions, his or her compensation would be $70,000 for the year. The manager is considering trying to boost his or her compensation for the year by increasing the firm's return on assets. The manager is thinking of increasing the return on assets by buying a new piece of equipment that would lower operating expenses (and cash outflows) by $X this year (before deduction of the added depreciation expense due to the purchase of the equipment). The equipment costs $70,000 and would be depreciated by $7,000 by the end of the year. What is the minimum that X must be in order for it to be worthwhile for the manager to purchase the equipment? Assume that the equipment is paid for with cash this year.
Answer #1

Currently, Firm's Return on Assets = Earnings / Total Asset at year end = 1,50,000 / 15,00,000 = 10%

For it to be worthwhile for the manager to purchase the equipment, the return on asset after the purchase must be at least 10%.

Now, Total Asset value at year end after equipment purchase = $ 15,00,000 + 70,000 (Equipment Added) - 7,000 (Depreciation) -70,000 (Cash Outflow for asset purchase) = $ 14,93,000.

Thus earnings required to maintain a Return on Asset of 10% on total asset of $ 14,93,000 = $ 14,93,000 * 10% = $ 1,49,300.

Current Earnings = $ 1,50,000.

Depreciation for equipment purchased = $ 7,000.

Earnings after depreciation = 1,50,000 - 7,000 = $ 1,43,000.

Thus, lowering of operating expenses required (i.e. 'X') = Required Earnings - Current Earnings after depreciation

= $ 1,49,300 - $ 1,43,000

= $ 6,300.

Hence, the minimum that X must be in order for it to be worthwhile for the manager to purchase the equipment is $ 6,300.